# On the estimate $M(x)=o(x)$ for Beurling generalized numbers

### v1

@@ Abstract.
Beurling generalized numbers; Beurling primes; mean-value vanishing of the Möbius function; prime number theorem; PNT equivalences
@@ 1. Introduction
(1.1) $M(x)=o(x),$
where $M$ stands for the sum function of the Möbius function, in the sense that they are deducible one another by quite simple real variable arguments (cf. [8, 14]). The counterparts of this and several other classical “PNT equivalences” for Beurling generalized numbers have been thoroughly investigated in the past decade [4, 6, 7, 10, 11]; see also the survey articles [9, 12]. It has been discovered that various of the different implications may fail without extra hypotheses, and conditions have also been determined under which the implications do or do not hold in the Beurling context.
@@
According to [4, Theorem 2.1], the PNT does imply (1.1) if one additionally assumes that the generalized integer counting function $N$ satisfies the bound $N(x)\ll x$. Therefore, if one would like to attempt to disprove the implication “PNT $\Rightarrow$ (1.1)”, one would need to seek counterexamples among those number systems for which $N(x)/x$ is unbounded. This observation in combination with [11, Proposition 6.1, p. 60] led Kahane to propose [^1] Orally communicated in April 2017 during an author’s visit to Jean-Pierre Kahane and Eric Saïas in Paris. the absolutely continuous prime measure [^2] Even if one would mainly be interested in discrete Beurling number systems, considering “continuous analogs” is a useful device for the construction of counterexamples in Beurling’s number theory, as they provide flexible templates that could usually be discretized by means of a number of available techniques (cf. [3, 5, 11]). So, in a broad sense [1, 11], a Beurling generalized number system is merely a pair of non-decreasing right continuous functions $N$ and $\Pi$ with $N(1)=1$ and $\Pi(1)=0$, both having support in $[1,\infty)$, and subject to the relation $\mathrm{d}N=\exp^{\ast}(\mathrm{d}\Pi)$. (we write $\log_{k}x=\log(\log_{k-1}x)$)
@@
as a possible continuous candidate to refute “PNT $\Rightarrow$ (1.1)”. In fact, we obviously have $\Pi_{\text{K}}(x)\sim x/\log x$, while the quoted result from [11] yields $N_{\text{K}}(x)/x\to\infty$ for its associated generalized integer distribution function, determined via the relation $\mathrm{d}N_{K}=\exp^{\ast}(\mathrm{d}\Pi_{\text{K}})$ where the exponential is taken with respect to the (multiplicative) convolution of measures [11]. The analog of the sum of the Möbius function in this case is $M_{K}$ with $\mathrm{d}M_{K}$ defined as the convolution inverse of $\mathrm{d}N_{\text{K}}$, or equivalently, $\mathrm{d}M_{\text{K}}=\exp^{\ast}(-\mathrm{d}\Pi_{\text{K}})$.
We have found here that $M_{K}(x)=o(x)$; consequently, the prime measure (1.2) fails to deliver a counterexample for “PNT $\Rightarrow$ (1.1)”. This follows at once from the following theorem, which is our main result. We mention that primitives of all measures considered in this article are right-continuous and supported on the interval $[1,\infty)$.
@@ 2. Preliminaries on convolution calculus
$\int_{1^{-}}^{x}\mathrm{d}A\ast\mathrm{d}B:=\iint_{1\leq uv\leq x}\mathrm{d}A(u)\mathrm{d}B(v)=\int_{1^{-}}^{x}A\left(\frac{x}{u}\right)\mathrm{d}B(u)=\int_{1^{-}}^{x}B\left(\frac{x}{u}\right)\mathrm{d}A(u).$
@@
We shall make use of the multiplication by $\log$ operator, namely, the operator $L$ acting on measures as
@@
The reader should think of (2.2) as a general form of Chebyshev’s classical identity from elementary prime number theory.
@@
$\int_{1^{-}}^{\infty}\frac{|\exp^{\ast}(\mathrm{d}R(u))|}{u}\leq\int_{1^{-}}^{\infty}\frac{\exp^{\ast}(|\mathrm{d}R(u)|)}{u}=\exp\left(\int_{1^{-}}^{\infty}\frac{|\mathrm{d}R(u)|)}{u}\right)<\infty.$
@@
$\int_{1^{-}}^{\infty}\frac{\mathrm{d}F^{+}(u)}{u^{\sigma}}=\exp\left(\int_{1^{-}}^{\infty}\frac{\mathrm{d}E(u)}{u^{\sigma}}\right)\ll\frac{1}{(\sigma-1)^{\varepsilon}}.$
An quick elementary Tauberian argument shows that
$\int_{1^{-}}^{x}\frac{\mathrm{d}F^{+}(u)}{u}\ll\log^{\varepsilon}x.$
Noticing that $\tilde{E}(x)=\int_{1^{-}}^{x}\log u\,\mathrm{d}E(u)=\log u\,E(u)-\int_{1}^{x}u^{-1}E(u)\,\mathrm{d}u=o(x)$ and employing property (2.2), we deduce that
$H^{+}(x)=\int_{1^{-}}^{x}\mathrm{d}\tilde{E}\ast\mathrm{d}F^{+}=\int_{1^{-}}^{x}\tilde{E}\left(\frac{x}{u}\right)\ast\mathrm{d}F^{+}(u)\ll x\int_{1^{-}}^{\infty}\frac{\mathrm{d}F^{+}(u)}{u}\ll x\log^{\varepsilon}x.$
@@
$\leq x\log^{\varepsilon}x+x\log^{\varepsilon}x\int_{1}^{x}\frac{\mathrm{d}u}{u\log^{a}(e+u)}\ll\log^{1-b}x.$
@@
has primitive $1/x$. Using this formula,
@@
$\int_{1^{-}}^{\infty}\frac{\mathrm{d}B^{+}(u)}{u^{\sigma}}=\exp\left(\int_{1^{-}}^{\infty}\frac{\mathrm{d}A^{+}(u)}{u^{\sigma}}\right)=b_{1}\log\left(\frac{1}{\sigma-1}\right)+b_{2}+o(1).$
@@
$\quad\int_{1^{-}}^{x}\frac{\mathrm{d}B(u)}{u}=b_{1}\log_{2}x+b_{2}+b_{1}\gamma+o(1)$
@@
(4.2) $B_{-}(x)=o\left(\frac{x}{\log x}\right).$

### v2

@@ Abstract.
Beurling generalized numbers; Beurling generalized primes; mean-value vanishing of the Möbius function; prime number theorem; PNT equivalences
@@ 1. Introduction
(1.1) $M(x)=o(x),\qquad x\to\infty,$
where $M$ stands for the sum function of the Möbius function, in the sense that they are deducible from one another by quite simple real variable arguments (cf. [8, 14]). The counterparts of this and several other classical “PNT equivalences” for Beurling generalized numbers have been thoroughly investigated in the past decade [4, 6, 7, 10, 11]; see also the survey articles [9, 12]. It has been discovered that various of the different implications may fail without extra hypotheses, and conditions have also been determined under which the implications do or do not hold in the Beurling context.
@@
According to [4, Theorem 2.1], the PNT does imply (1.1) if one additionally assumes that the generalized integer counting function $N$ satisfies the bound $N(x)\ll x$. Therefore, if one would like to attempt to disprove the implication “PNT $\Rightarrow$ (1.1)”, one would need to seek counterexamples among those number systems for which $N(x)/x$ is unbounded. This observation in combination with [11, Proposition 6.1, p. 60] led Kahane to propose [^1] Orally communicated in April 2017 during an author’s visit to Jean-Pierre Kahane and Eric Saïas in Paris. the absolutely continuous prime measure [^2] Even if one would mainly be interested in discrete Beurling number systems, considering “continuous analogs” is a useful device for the construction of counterexamples in Beurling number theory, as they provide flexible templates that could usually be discretized by means of a number of available techniques (cf. [3, 5, 11]). So, in a broad sense [1, 11], a Beurling generalized number system is merely a pair of non-decreasing right continuous functions $N$ and $\Pi$ with $N(1)=1$ and $\Pi(1)=0$, both having support in $[1,\infty)$, and subject to the relation $\mathrm{d}N=\exp^{\ast}(\mathrm{d}\Pi)$. (we write $\log_{k}x=\log(\log_{k-1}x)$)
@@
as a possible candidate to refute “PNT $\Rightarrow$ (1.1)”. In fact, we obviously have $\Pi_{\text{K}}(x)\sim x/\log x$, while the quoted result from [11] yields $N_{\text{K}}(x)/x\to\infty$ for its associated generalized integer distribution function, determined via the relation $\mathrm{d}N_{K}=\exp^{\ast}(\mathrm{d}\Pi_{\text{K}})$ where the exponential is taken with respect to the (multiplicative) convolution of measures [11]. The analog of the sum of the Möbius function in this case is $M_{K}$ with $\mathrm{d}M_{K}$ defined as the convolution inverse of $\mathrm{d}N_{\text{K}}$, or equivalently, $\mathrm{d}M_{\text{K}}=\exp^{\ast}(-\mathrm{d}\Pi_{\text{K}})$.
We have found that $M_{K}(x)=o(x)$; consequently, the prime measure (1.2) fails to deliver a counterexample for “PNT $\Rightarrow$ (1.1)”. This follows at once from the following theorem, which is our main result. We mention that primitives of all measures considered in this article are right-continuous and supported on the interval $[1,\infty)$.
@@
The author thanks Frederik Broucke, Gregory Debruyne, and Harold G. Diamond for their useful comments.
@@ 2. Preliminaries on convolution calculus
$\int_{1^{-}}^{x}\mathrm{d}A\ast\mathrm{d}B=\iint_{1\leq uv\leq x}\mathrm{d}A(u)\mathrm{d}B(v)=\int_{1^{-}}^{x}A\left(\frac{x}{u}\right)\mathrm{d}B(u)=\int_{1^{-}}^{x}B\left(\frac{x}{u}\right)\mathrm{d}A(u).$
@@
We shall make use of the operator $L$ of multiplication by $\log$, namely,
@@
The reader should think of (2.2) as a general form of Chebyshev’s classical identity from elementary prime number theory [^3] That is, the identity $\sum_{n\leq x}\log n=\sum_{n\leq x}\psi(x/n)$, where $\psi$ is the Chebyshev function..
@@
$\int_{1^{-}}^{\infty}\frac{|\exp^{\ast}(\mathrm{d}R(u))|}{u}\leq\int_{1^{-}}^{\infty}\frac{\exp^{\ast}(|\mathrm{d}R(u)|)}{u}=\exp\left(\int_{1^{-}}^{\infty}\frac{|\mathrm{d}R(u)|}{u}\right)<\infty.$
@@
(3.4) $\int_{1^{-}}^{\infty}\frac{\mathrm{d}F^{+}(u)}{u^{\sigma}}=\exp\left(\int_{1^{-}}^{\infty}\frac{\mathrm{d}E(u)}{u^{\sigma}}\right)\ll\frac{1}{(\sigma-1)^{\varepsilon}}\>.$
An quick elementary Tauberian argument shows that $\int_{1^{-}}^{x}u^{-1}\mathrm{d}F^{+}(u)\ll\log^{\varepsilon}x.$ Indeed, taking $\sigma=1+1/\log x$ in (3.4), we obtain
$\int_{1^{-}}^{x}\frac{\mathrm{d}F^{+}(u)}{u}\leq e\int_{1^{-}}^{\infty}\exp\left(-\frac{\log u}{\log x}\right)\frac{\mathrm{d}F^{+}(u)}{u}\ll\log^{\varepsilon}x.$
Noticing now that $\tilde{E}(x)=\int_{1^{-}}^{x}\log u\,\mathrm{d}E(u)=E(x)\log x-\int_{1}^{x}u^{-1}E(u)\,\mathrm{d}u=o(x)$ and employing property (2.2), we deduce that
$H^{+}(x)=\int_{1^{-}}^{x}\mathrm{d}\tilde{E}\ast\mathrm{d}F^{+}=\int_{1^{-}}^{x}\tilde{E}\left(\frac{x}{u}\right)\mathrm{d}F^{+}(u)\ll x\int_{1^{-}}^{\infty}\frac{\mathrm{d}F^{+}(u)}{u}\ll x\log^{\varepsilon}x.$
@@
$\leq x\log^{\varepsilon}x+x\log^{\varepsilon}x\int_{1}^{x}\frac{\mathrm{d}u}{u\log^{a}(e+u)}\ll x\log^{1-b}x.$
@@
has primitive $1/x$. Using this formula, one obtains
@@
$\int_{1^{-}}^{\infty}\frac{\mathrm{d}B(u)}{u^{\sigma}}=\int_{1^{-}}^{\infty}\frac{\mathrm{d}B^{+}(u)}{u^{\sigma}}+o(1)=b_{1}\log\left(\frac{1}{\sigma-1}\right)+b_{2}+o(1).$
@@
$\quad\int_{1^{-}}^{x}\frac{\mathrm{d}B(u)}{u}=b_{1}\log_{2}x+b_{2}+b_{1}\gamma+o(1),$
@@
(4.2) $B^{-}(x)=o\left(\frac{x}{\log x}\right).$